\section*{References}

\begin{thebibliography}{99}
\bibitem{Szmy97}
	 R. Szmytkowski,
   Journal of Physics B 30 (1997) 825  
   [erratum: Journal of Physics B 30 (1997) 2747;
   addendum: arXiv:physics/9902050].
\bibitem{Szmy02b}
   R. Szmytkowski,
   Journal of Physics B 35 (2002) 1379. 
\bibitem{Szmy04}
   R. Szmytkowski, K. Mielewczyk,
   Journal of Physics B 37 (2004) 3961. 
\bibitem{Szmy02a}
   R. Szmytkowski,
   Physical Review A 65 (2002) 012503. 
\bibitem{Miel06}
   K. Mielewczyk, R. Szmytkowski,
   Physical Review A 73 (2006) 022511  
   [erratum:  Physical Review A 73 (2006) 039908].
\bibitem{Szmy11}
   R. Szmytkowski, P. Stefa{\'n}ska,
   e-print arXiv:1102.1811.
\bibitem{Stef12}
   P. Stefa{\'n}ska, R. Szmytkowski,
   International Journal of Quantum Chemistry 112 (2012) 1363. 
\bibitem{Szmy12}
   R. Szmytkowski, P. Stefa{\'n}ska,
   Physical Review A 85 (2012) 042502. 
\bibitem{Szmy14}
   R. Szmytkowski and P. Stefa{\'n}ska, 
   Physical Review A 89 (2014) 012501.	
\bibitem{Stef15}
	P. Stefa{\'n}ska, 
	Physical Review A 92 (2015) 032504.
\bibitem{Mohr14}
	CODATA Internationally recommended 2014 values of the fundamental physical constants,
	http://physics.nist.gov/cuu/Constants/index.html.
\bibitem{Rutk07} 
	A. Rutkowski, A. Poszwa, 
	Physical Review A 75 (2007) 033402.
	

\end{thebibliography}

\documentclass[reviewcopy]{elsarticle}

\usepackage[reviewcopy]{adndt}
\usepackage{longtable}

\usepackage{amsmath}

\usepackage{amssymb}

\usepackage{lscape}

\biboptions{square,sort&compress}
\bibpunct[]{[}{]}{,}{n}{}{;}
\begin{document}

\begin{frontmatter}

\journal{Atomic Data and Nuclear Data Tables}


\title{Magnetizabilities of relativistic hydrogenlike atoms \\ in some arbitrary discrete energy eigenstates}

  \author{Patrycja Stefa{\'n}ska\corref{cor1}}
  \ead{pstefanska@mif.pg.gda.pl}

  \cortext[cor1]{Corresponding author.}

  \address{Atomic Physics Division,
Department of Atomic, Molecular and Optical Physics, 
Faculty of Applied Physics and Mathematics, \\
Gda{\'n}sk University of Technology, 
Narutowicza 11/12, 80--233 Gda{\'n}sk, Poland}

\date{\today}

\begin{abstract}  
We present the results of numerical calculations of magnetizability ($\chi$) of the relativistic one-electron atoms with a pointlike, spinless and motionless nuclei of charge $Ze$. Exploiting the analytical formula for $\chi$ recently derived by us [P. Stefa{\'n}ska, 2015], valid for an arbitrary discrete energy eigenstate, we have found the values of the magnetizability for the ground state and for the first and the second set of excited states (i.e.: $2s_{1/2}$, $2p_{1/2}$, $2p_{3/2}$, $3s_{1/2}$, $3p_{1/2}$, $3p_{3/2}$, $3d_{3/2}$, and $3d_{5/2}$) of the Dirac one-electron atom. The results for ions with the atomic number  $1 \leqslant Z \leqslant 137$ are given in 14 tables. The  comparison of the numerical values of magnetizabilities for the ground state and for each states belonging to the first set of excited states of selected hydrogenlike ions, obtained with the use of two different values of the fine-structure constant, i.e.: $\alpha^{-1}=137.035 \: 999 \: 139$ (CODATA 2014) and $\alpha^{-1}=137.035 \: 999 \: 074$ (CODATA 2010), is also presented.\\

\noindent
\textbf{Keywords:} Hydrogenlike atom, Magnetizability, Electromagnetic moments, Dipole moment, Magnetic field.\\

\begin{center}
\large{\textbf{Published as: At. Data Nucl. Data Tables 108 (2016) 193--210}
\\*[1ex]
\textbf{doi: 10.1016/j.adt.2015.09.001}\\*[5ex]}
\end{center}
\end{abstract}

\end{frontmatter}




\newpage

\tableofcontents
\listofDtables
\listofDfigures
\vskip5pc


\section{Introduction}

Interaction of atoms and molecules with electromagnetic field is undoubtedly one of the most commonly reported physical processes, both theoretically and experimentally. For the simplest systems, like one-electron atoms, there are some analytical methods of calculating many atomic parameters, such as the polarizability or the magnetizability. One of such useful tool is the Sturmian expansion of the first-order generalized Dirac--Coulomb Green function \cite{Szmy97}, proposed by Szmytkowski in 1997. In the series of papers \cite{Szmy02b,Szmy04,Szmy02a,Miel06,Szmy11,Stef12,Szmy12,Szmy14} published by his group over the period of past several years, it has been used in perturbation-theory calculations of some electromagnetic properties of the relativistic hydrogenlike atoms in the ground state, with a poinlike, spinless and motionless nuclei of charge $Z e$.

Recently, we have shown that the usefulness of this method goes beyond the study of the atomic ground state. In Ref.\ \cite{Stef15} we derived analytically an expression for the magnetizability of the Dirac one-electron atom (with regard to its nucleus we impose the same assumptions as above) in an arbitrary discrete energy eigenstate, characterized by the set of quantum numbers $\{n, \kappa, \mu \}$, in which $n$ denotes the radial quantum number, the Dirac quantum number $\kappa$ is an integer different form zero, whereas $\mu=-|\kappa|+\frac{1}{2}, -|\kappa|+\frac{3}{2}, \ldots, |\kappa|-\frac{1}{2}$ is the magnetic quantum number. The final result has the following form:
\begin{eqnarray}
\chi \equiv \chi_{n\kappa\mu}&=&\frac{\alpha^2 a_0^3}{Z^2}\frac{1}{128(4\kappa^2-1)^2 N_{n \kappa}}
\Bigg\{
\Theta_{n\kappa\mu}^{(\textrm{I})}+\sum_{\kappa'}
\frac{\eta_{\kappa\mu}^{(+)}\delta_{\kappa',-\kappa+1}+\eta_{\kappa\mu}^{(-)} \delta_{\kappa',-\kappa-1}}{N_{n\kappa}+\kappa'}
\nonumber \\
&&\times
\Bigg[
\Theta_{n\kappa}^{(\textrm{II})}+\frac{n!(n^2+2n\gamma_{\kappa}+\kappa^2)\Gamma(n+2\gamma_{\kappa}+1)}{(N_{n\kappa}-\kappa)(\gamma_{\kappa'}-\gamma_{\kappa}-n+1)\Gamma(2\gamma_{\kappa'}+1)}\sum_{k=0}^n\sum_{p=0}^n \widetilde{\mathcal{Z}}_{\kappa \kappa'}^{n}(k)\widetilde{\mathcal{Z}}_{\kappa \kappa'}^{n}(p)
\nonumber \\
&&\quad 
\times
{}_3F_2 
\left(
\begin{array}{c} 
\gamma_{\kappa'}-\gamma_{\kappa}-k-1,\: 
\gamma_{\kappa'}-\gamma_{\kappa}-p-1,\: 
\gamma_{\kappa'}-\gamma_{\kappa}-n+1 \\
\gamma_{\kappa'}-\gamma_{\kappa}-n+2,\:
2\gamma_{\kappa'}+1
\end{array}
;1 
\right) 
\Bigg]
\Bigg\},
\label{eq:1}
\end{eqnarray}
where $\alpha$ is the Sommerfeld's fine structure constant, $\Gamma(\zeta)$ denotes the Euler's gamma function, ${}_3F_2$ is the generalized hypergeometric function, while
\begin{equation}
\eta_{\kappa\mu}^{(\pm)}=(4\kappa^2-1)^2-4\mu^2(2\kappa\pm1)^2,
\label{eq:2}
\end{equation} 
\begin{equation}
\Theta_{n\kappa\mu}^{(\textrm{I})}=-256\kappa^2 \mu^2
\left[
2\kappa^2(n+\gamma_{\kappa})^3+(n+\gamma_{\kappa})(5n^2+10n \gamma_{\kappa}+2\gamma_{\kappa}^2-2\kappa^2+1)N_{n\kappa}^2-\kappa(3n^2+6n \gamma_{\kappa}+4\gamma_{\kappa}^2-\kappa^2)N_{n\kappa}
\right],
\label{eq:3}
\end{equation}
\begin{equation}
\Theta_{n\kappa}^{(\textrm{II})}=2(2n+2\gamma_{\kappa}+1)(\kappa-N_{n \kappa})N_{n\kappa}^2
\left[
5(n+\gamma_{\kappa})(n+\gamma_{\kappa}+1)-3(\gamma_{\kappa}^2-1)
\right],
\label{eq:4}
\end{equation}
\begin{equation}
\widetilde{\mathcal{Z}}_{\kappa \kappa'}^{n}(k)= \frac{(-)^{k} \left[2(N_{n \kappa}-\kappa)+(n-k)(\kappa+\kappa')\right]}{k!(n-k)!} \:
 \frac{\Gamma(\gamma_{\kappa}+\gamma_{\kappa'}+k+2)}{\Gamma(k+2\gamma_{\kappa}+1)}
\label{eq:5}
\end{equation}
and analogously for $\widetilde{\mathcal{Z}}_{\kappa \kappa'}^{n}(p)$, with
\begin{equation}
N_{n\kappa}=\sqrt{n^2+2n\gamma_{\kappa}+\kappa^2},
\label{eq:6}
\end{equation}
and
\begin{equation}
\gamma_{\kappa}=\sqrt{\kappa^2-(\alpha Z)^2}.
\label{eq:7}
\end{equation}

The above result has been exhaustively verified by us, both analytically and numerically. In Ref.\ \cite{Stef15} we have shown that it remains valid for {{\em an arbitrary} discrete energy eigenstate. However, the aforementioned article contains only two representational tables with values of the relativistic magnetizabilities for some excited states of selected hydrogenlike ions. In this work, we present a more comprehensive numerical data in the form of 14 tables comprising the results for the atomic ground state $1s_{1/2}$ and for each state belonging to the first and second set of excited states, i.e.: $2s_{1/2}$, $2p_{1/2}$, $2p_{3/2}$, $3s_{1/2}$, $3p_{1/2}$, $3p_{3/2}$, $3d_{3/2}$ and $3d_{5/2}$, having regard all possible values of the magnetic quantum number. 

Present calculations have been performed with the current value $137.035 \: 999 \: 139$ of the inverse of the fine-structure constant recommended by the Committee on Data for Science and Technology (CODATA) \cite{Mohr14}, in contrast to those described in Ref.\ \cite{Stef15}, in which we have used $\alpha^{-1}=137.035 \: 98 \:95$ (CODATA 1986), to be able to compare our results with the previous results of other authors \cite{Rutk07}. To show how the change in the value of the fine structure constant affects the value of the magnetizability, additionally, we performed calculations for $\chi$ with $\alpha^{-1}=137.035 \: 999 \: 074$, recommended by CODATA 2010 (i.e. immediately before the currently valid value of this constant) and compared them with the corresponding values from Tables \ref{tab:1s_1-2} -- \ref{tab:2p_3-2_mu_3-2}. The appropriate juxtaposition of numerical values of $\chi$ for the ground state and the first excited states of selected hydrogenlike ions are shown in Table \ref{tab:comparison}.



\ack
I wish to thank Professor R.\ Szmytkowski for many stimulating discussions, especially for every advice he gave me during the preparation of this work.



\clearpage

\newpage

\TableExplanation

In all the tables we have used the following notation:

\bigskip
\renewcommand{\arraystretch}{1.0}


\end{landscape}

\end{document}